OpenStudy (anonymous):

Solve for x: x+1=5/2g

OpenStudy (dangerousjesse):

Solve for x: \[5 g = 2 (x+1) \]Reverse the equality in \(5 g = 2 (x+1)\) in order to isolate x to the left hand side. \(5 g = 2 (x+1)\) is equivalent to \(2 (x+1) = 5 g:\) \[2 (x+1) = 5 g \]Divide both sides by a constant to simplify the equation. Divide both sides by 2: \[x+1 = \frac{5 g}{2}\]Solve for x. Subtract 1 from both sides:

OpenStudy (anonymous):

@DangerousJesse Thanks.